OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (strawberry17):

is the solution set supposed to be: ? y = x - 5 y = -x -4

OpenStudy (anonymous):

oh sorry they both have <=

OpenStudy (strawberry17):

could you rewrite each with the correct signs?

OpenStudy (strawberry17):

@src4235

OpenStudy (anonymous):

yes, sorry y<=x-5 y<=-x-4

OpenStudy (strawberry17):

how can they have both an equals sign and an inequality sign in one equation??

OpenStudy (anonymous):

that's an make shift less than or equal to sign

OpenStudy (strawberry17):

oh is it supposed to be ≤ on both? :)

OpenStudy (anonymous):

yes that is correct:) lol

OpenStudy (strawberry17):

okay, I will solve :)

OpenStudy (anonymous):

thank you!:)

OpenStudy (strawberry17):

I got (10, 1) and (1, -10). (10, 1) fits the first inequality and (1, -10) fits the second inequality but not vise versa. So it must be both points?

OpenStudy (anonymous):

yes thank you so much!:)

OpenStudy (strawberry17):

ur welcome! :)